# Feel-Good Thompson Sampling for Contextual Bandits: a Markov Chain Monte Carlo Showdown

### v1

@@ Abstract
Thompson Sampling (TS) is widely used to address the exploration/exploitation tradeoff in contextual bandits, yet recent theory shows that it does not explore aggressively enough in high-dimensional problems. Feel-Good Thompson Sampling (FG-TS) addresses this by adding an optimism bonus that biases toward high-reward models, and it achieves the asymptotically minimax-optimal regret in the linear setting when posteriors are exact. However, its performance with approximate posteriors—common in large-scale or neural problems—has not been benchmarked. We provide the first systematic study of FG-TS and its smoothed variant (SFG-TS) across eleven real-world and synthetic benchmarks. To evaluate their robustness, we compare performance across settings with exact posteriors (linear and logistic bandits) to approximate regimes produced by fast but coarse stochastic-gradient samplers. Ablations over preconditioning, bonus scale, and prior strength reveal a trade-off: larger bonuses help when posterior samples are accurate, but hurt when sampling noise dominates. FG-TS generally outperforms vanilla TS in linear and logistic bandits, but tends to be weaker in neural bandits. Nevertheless, because FG-TS and its variants are competitive and easy-to-use, we recommend them as baselines in modern contextual-bandit benchmarks. Finally, we provide source code for all our experiments in https://github.com/SarahLiaw/ctx-bandits-mcmc-showdown.
**footnotetext: These authors contributed equally to this work.$\dagger$\$$\dagger$\$footnotetext: Work done while an intern at Windsurf and Cognition.
@@ 1 Introduction
The stochastic multi-armed bandit model (Berry and Fristedt, 1985; Bubeck and Cesa-Bianchi, 2012; Lattimore and Szepesvári, 2020) is one of the most important frameworks for sequential decision making under uncertainty. Among its variants, the contextual bandit (Langford and Zhang, 2007) has enabled vast breakthroughs in numerous real-world applications such as recommendation (Yue et al., 2012), predictive control (Lin et al., 2023a), healthcare (Durand et al., 2018), and prompt optimization (Dwaracherla et al., 2024). In each round, the contextual bandit observes a $d$-dimensional feature vector (the “context”) for each of its arms, pulls one of them, and receives a reward. The bandit’s goal is to choose actions to maximize the total reward over a finite horizon $T$. This limited bandit feedback highlights the exploration/exploitation dilemma: should one choose the myopically better arm to maximize an immediate reward, or an under-sampled arm to potentially improve future rewards?
@@
[…]
@@
    Neural contextual bandits. We study the performance of FG-TS and SFG-TS with TS in six real-world datasets, through the Langevin Monte Carlo (LMC) MCMC algorithm. In the neural setting, the additional computations done by more sophisticated MCMC algorithms render the computations too expensive to be practical. Nevertheless, we compare the performance of these methods with well-studied algorithms, such as Neural-$\epsilon$-Greedy and Neural UCB.
Our experiments demonstrate that while FG-TS performs the best (through the LMC algorithm) in the linear contextual bandit setting, and while SFG-TS performs the best (through the MALA algorithm) in the logistic contextual bandit setting, FG-TS (and its smoothed variants) tends to be weaker in the neural contextual bandit setting. Perhaps surprisingly, our experiments demonstrate several instances where the performance of FG-TS degenerates very rapidly in the neural setting, where TS remains far more competitive. Since FG-TS favors aggressive exploration in contextual bandits, it is expected to be more amenable for real-world datasets; however, our experiments indicate that this is not the case.
@@ 3 Algorithms
This section describes the different algorithmic design principles we considered in our simulations in Section 4. These algorithms include (Neural-)$\epsilon$-Greedy, Linear UCB, LinTS, Neural UCB, and our main focus: Markov Chain Monte Carlo methods, with its many variants.
@@ 3.1 Markov Chain Monte Carlo Methods
[…]
@@
$\widetilde{\mu}=\nabla U(\widetilde{\theta})=\nabla U_{\mathrm{data}}(\widetilde{\theta})$ $+2\alpha\|\widetilde{\theta}\|,\quad\quad g_{k}(\theta_{t})=\nabla U_{\mathrm{data}}^{(\mathcal{B}_{k})}(\theta_{t})-\nabla U_{\mathrm{data}}^{(\mathcal{B}_{k})}(\widetilde{\theta})+\widetilde{\mu},$ (3)
@@ 4 Empirical Evaluation
In this section, we present an empirical evaluation on two simulated bandit generators (linear, logistic) and six real-world classification tasks. All our experiments are run on AWS EC2 servers with P3/G4 GPUs and 6 vCPUs, Google Cloud Compute servers with L4 GPUs, and N2 instances with 8 vCPUs.
@@ 4.2 Testbeds, Baselines, and Common Experimental Setup
[…]
@@
Real-World Tasks. Following Riquelme et al. (2018), we evaluate the MCMC Thompson Sampling algorithms on six UCI datasets—Adult, Shuttle, MagicTelescope, Mushroom, Jester, and Covertype, along with a realistic Financial dataset. Finally, we also evaluate the algorithms with the MNIST_784 image set (Krizhevsky, 2009; Elmachtoub et al., 2017). These are standard datasets studied in contextual bandits and multi-agent settings (Riquelme et al., 2018; Chaudhari et al., 2024), as they demonstrate a variety of features such as stochasticity of rewards, size, number of optimal actions, etc. Following the protocol of Riquelme et al. (2018); Kveton et al. (2020), each $N$-class example $x\!\in\!\mathbb{R}^{d}$ is converted into $K$ arm–context vectors $\smash{x^{(1)},\dots,x^{(k)}}$, where $x^{(i)}=x\cdot e_{i}$ where $e_{i}$ is the $i$’th standard basis vector of $\mathbb{R}^{d}$. Pulling arm $j$ returns reward $1$ if $j$ matches the ground-truth label, and 00 otherwise. We provide statistics of the benchmark datasets in Appendix F.
Network Architectures and Training Protocol. We adopt the training protocol and hyperparameter search strategy of Xu et al. (2022b): each method chooses the better of a fully connected two-layer MLP (width $100$) and four-layer MLP (width $50$); the activation is selected from {ReLU, LeakyReLU}. Networks are updated for $100$ SGD steps every round. FG-augmented variants inherit identical network architectures, optimizers, and training schedules as their non-FG counterparts to ensure fair comparisons. Unless stated otherwise, we run each bandit for $T=10{,}000$ rounds (Covertype: $T=15{,}000$). Additional implementation minutiae appear in Appendix A of Riquelme et al. (2018).
@@ 5 Results
Linear contextual bandits. Table 1 (first 3 columns) reports the final cumulative regret (mean $\pm$ sample standard deviation over 10 seeds) in the synthetic linear bandit setting. Across both dimensions $d\in\{20,40\}$, MALATS consistently attains the lowest cumulative regret: $<62$ when $d=20$ (vs 62.6 for unadjusted LMCTS, 73.0 for LinUCB when $\beta=10^{3}$) and around 100.6 when $d=40$ (vs 129.1 for LMCTS, 126.3 for LinUCB). Adding a feel-good optimism bonus (FGMALATS and SFGMALATS) of $\lambda=0.5$ does not improve regret in these Gaussian-conjugate settings. However, for smaller values of $\lambda$, such as $\lambda=0.01$ in Table 4, the performance of SFGMALATS improves significantly to 56.2 $\pm$ 22.8, surpassing the other vanilla-TS variants. When $d=40$, PHMCTS incurs $162.2\pm 12.5$ and SVRGLMCTS incurs 19236.8 $\pm$ 15560.8, indicating that shifting the posterior due to preconditioning and variance-reduction distorts the quadratic curvature of the true posterior.
[…]
Neural contextual bandits. In Table 2, we report the final cumulative regret of a subset of our algorithms on six real-world problems, deferring the full experimental result to Section E.5. We also evaluate the quality of the learned policy via simple regret in Table 3. When the posterior is approximated by SFG-TS, we observe a partial reversal: LMCTS becomes competitive with, and often outperforms, its feel-good extensions on five of the six datasets. The only advantage of the feel-good bonus appears in Magic Telescope, where SFGLMCTS reduces cumulative regret by $29$ and achieves the lowest simple regret at $82.8\pm 8.9$. Meanwhile, Neural-$\epsilon$-Greedy and NeuralUCB remain competitive, achieving the lowest regret on Mushroom and Adult, respectively. Finally, simple regret is lowest for the greedy neural methods on three of the six tasks.
@@ 5.1 Ablation Studies
(f) MNIST
Figure 1: Mean regret comparison on simulated bandit problems. The shaded band around each mean curve represents ±$\pm 1$ sample standard deviation across 5 independent runs.
Table 4: Final cumulative regret for synthetic datasets. Sweeps over feel-good parameter $\lambda$, at fixed $\beta=10^{3}$.
@@ 6 Discussion
FG in Neural Contextual Bandits. Despite the stronger theoretical guarantees of FG-TS in the linear setting, the performance of FG-TS in the neural setting tends to rapidly degenerate, rendering vanilla-TS more competitive. As aforementioned, stochastic gradients already inject noise that provide large amounts of “free” exploration; here, the FG bonus may push the sampler further away from high-density regions. Moreover, neural posteriors are highly non-Gaussian and poorly captured by the quadratic surrogates used in the FG derivation; hence, optimism may increase modeling errors and lead to sub-optimal arm choices. Thus, vanilla LMCTS remains a strong baseline on most of the real world tasks. Corroborating this, our experiments show that in neural bandits, moderate exploration—be it through SGD noise or $\epsilon$-greedy perturbations—beats the optimism induced by FG-TS.
Preconditioning. Our preconditioned samplers perform Langevin updates in a whitened parameter space by injecting Gaussian noise with covariance proportional to the inverse design matrix $\mathbf{V}_{t}^{-1}$. When $\mathbf{V}_{t}$ is ill-conditioned, as it typically is in high dimensions, some directions receive huge stochastic kicks while others hardly move, leading to erratic exploration and exploding variance, as visible in SVRGLMCTS’s large regret. These imply a necessity to regularize or clip the spectrum of $\mathbf{V}_{t}^{-1}$ to keep noise injection balanced across directions, and to also include a Metropolis filter when using a mass-matrix preconditioner. Finally, our results on PSFGHMCTS indicate that it has a higher standard deviation than mean. This skew serves as an argument for incorporating additional metrics (such as the median) for studying the aggregate performance of MCMC contextual bandit algorithms.
@@
Damping. We provide results for the LMC algorithms with damping for the linear contextual bandit setting in Table 29. These results show that overdamped (vanilla) Langevin methods significantly outperform any of the corresponding underdamped experiments, and the gap diminishes monotonically with the damping coefficient $\gamma$. Despite the strong theoretical guarantees of underdamped LMC, the performance of ULMC (and its variants) rapidly degenerates due to the mismatch of the approximated posteriors amplified by the magnitude of the damping.

### v3

@@ Abstract
Thompson Sampling (TS) is widely used to address the exploration/exploitation tradeoff in contextual bandits, yet recent theory shows that it does not explore aggressively enough in high-dimensional problems. Feel-Good Thompson Sampling (FG-TS) addresses this by adding an optimism bonus that biases toward high-reward models, and it achieves the asymptotically minimax-optimal regret in the linear setting when posteriors are exact. However, its performance with approximate posteriors, common in large-scale or neural problems, has not been benchmarked. We provide the first systematic study of FG-TS and its smoothed variant (SFG-TS) across fourteen real-world and synthetic benchmarks. To evaluate their robustness, we compare performance across settings with exact posteriors (linear and logistic bandits) to approximate regimes produced by fast but coarse stochastic-gradient samplers. Ablations over preconditioning, bonus scale, and prior strength reveal a trade-off: larger bonuses help when posterior samples are accurate, but hurt when sampling noise dominates. FG-TS generally outperforms vanilla TS in linear and logistic bandits, but tends to be weaker in neural bandits. Nevertheless, because FG-TS and its variants are competitive and easy-to-use, we recommend them as baselines in modern contextual-bandit benchmarks. Finally, we provide source code for all our experiments in https://github.com/SarahLiaw/ctx-bandits-mcmc-showdown.
[…]
@@ 1 Introduction
The stochastic multi-armed bandit model (Berry and Fristedt, 1985; Bubeck and Cesa-Bianchi, 2012; Lattimore and Szepesvári, 2020) is one of the most prevalent frameworks for sequential decision making under uncertainty. Among its variants, the contextual bandit (Langford and Zhang, 2007) has enabled vast breakthroughs in numerous real-world applications such as recommendation (Yue et al., 2012), predictive control (Lin et al., 2023a), healthcare (Durand et al., 2018), and prompt optimization (Dwaracherla et al., 2024). In each round, the contextual bandit observes a $d$-dimensional feature vector (the “context”) for each of its arms, pulls one of them, and receives a reward. The bandit’s goal is to choose actions to maximize the total reward over a finite horizon $T$. This limited bandit feedback highlights the exploration/exploitation dilemma: should one choose the myopically better arm to maximize an immediate reward, or an under-sampled arm to potentially improve future rewards?
@@
In this paper, we develop a benchmark for exploration methods, and provide extensions to deep neural networks. We examine a wide class of exploration algorithms that have been studied for the past two decades (Bubeck et al., 2009; Mnih et al., 2015; Riquelme et al., 2018; Kveton et al., 2020; Xu et al., 2022a; Lin et al., 2024; Anand et al., 2025b; Lin et al., 2023b): we compare a variety of well-established and recent posterior sampling algorithms for approximating the posterior distribution, under the lens of FG-TS and its smoothed variants for contextual bandits. This includes a novel optimal-regret Hamiltonian Monte-Carlo algorithm for SFG-TS, building on the works of Zhang (2022); Huix et al. (2023). We also provide a theoretical analyses of the runtime complexities, convergence guarantees, and regret bounds in Appendix D. Finally, all code and implementations to reproduce the experiments will be available open-source to provide a reproducible benchmark for future development. [^1] Our PyTorch implementation is available here and our pip installation instructions are available here.
@@
    Neural contextual bandits. We study the performance of FG-TS and SFG-TS with TS in real-world datasets, through the Langevin Monte Carlo (LMC) MCMC algorithm. In the neural setting, the additional computations done by more sophisticated MCMC algorithms render the computations too expensive to be practical. Nevertheless, we compare the performance of these methods with well-studied algorithms, such as Neural-$\epsilon$-Greedy and Neural UCB.
Our experiments show that FG-TS performs best (through the LMC algorithm) in the linear contextual bandit setting, and that SFG-TS performs best (through the MALA algorithm) in the logistic setting. However, FG-TS (and its smoothed variants) tends to be weaker in the neural contextual bandit setting. Perhaps surprisingly, we observe several instances where its performance degenerates rapidly, where TS remains far more competitive. A priori, one might expect FG-TS to be more amenable to real-world datasets because it favors more aggressive exploration (e.g., by escaping local optima or handling sparse or heterogeneous contexts); however, our experiments indicate that this is not the case.
@@ 3 Algorithms
[…]
@@ 3.1 Markov Chain Monte Carlo Methods
Langevin Monte Carlo (LMC). Xu et al. (2022b) proposed a Langevin MCMC to approximate the unknown posterior distribution of the parameter $\mathbf{\theta}^{*}$ upto a high precision: given a schedule of step-sizes $\{\eta_{t}\}_{t\geq 1}$ and inverse temperatures $\{\beta_{t}\}_{t\geq 1}$, LMC [^2] The Euler-Maruyama discretization of the Langevin SDE: $\smash{\mathrm{d}\theta_{s}=-\nabla\mathcal{L}_{t}(\theta_{s})+\sqrt{2\beta_{t}^{-1}}\mathrm{d}B(s)}$ samples a standard normal vector $\epsilon_{t,k}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and lets $\smash{\theta_{t,k+1}=\theta_{t,k}-\eta_{t}\nabla\mathcal{L}_{t}(\theta_{t,k-1})+({2\eta_{t}\beta_{t}^{-1}})^{1/2}\epsilon_{t,k}}$. For sufficiently large epoch lengths $K_{t}$, this approximately converges to the posterior distribution $\pi_{t}(\theta)\propto e^{-\beta_{t}L_{t}(\theta)}$}.
@@
[…]
@@ 4 Empirical Evaluation
In this section, we present an empirical evaluation on two simulated bandit generators (linear and logistic) and a suite of real-world classification tasks. Experiments are conducted on AWS EC2 servers with P3/G4 GPUs (6 vCPUs), Google Cloud Compute servers with L4 GPUs and N2 instances (8 vCPUs), and NVIDIA A100 GPU.
@@ 4.2 Testbeds, Baselines, and Common Experimental Setup
Linear Contextual Bandit. To assess the performance in a setting with a tractable true posterior, we use a linear contextual bandit environment. At each round $t\in[T]$, the agent observes a context $\smash{\mathcal{X}_{t}\sim\mathcal{N}(\mathbf{0}_{4},\mathbf{I}_{4})}$, chooses an action $x_{t}\in[K]$ with $K=5$, and receives a noisy linear reward $\smash{r_{t}=\phi(\mathcal{X}_{t},x_{t})^{\top}\theta^{\star}+\varepsilon_{t}}$, where $\smash{\varepsilon_{t}\sim\mathcal{N}(0,\sigma^{2}),\sigma=0.5}$ and $\smash{\theta^{\star}\in\mathbb{R}^{20}}$. We place a Gaussian prior, $\smash{\theta_{0}\sim\mathcal{N}\bigl(\mathbf{0}_{20},\sigma_{0}^{2}\mathbf{I}\bigr)}$ with $\sigma_{0}=0.01$. The feature map $\phi:\mathbb{R}^{4}\times[K]\to\mathbb{R}^{20}$ is the standard block concatenation $(\phi(\mathcal{X}_{t},0),\dots,\phi(\mathcal{X}_{t},K))$ where $\smash{\phi(\mathcal{X}_{t},i)=e_{i}\cdot\mathcal{X}_{t}}$ (where $e_{i}$ is the $i$’th standard basis vector). Now, $\theta^{\star}$ naturally decomposes into $k$ context-specific blocks. The time horizon is T=10  [^0] T=10\,000. Since the likelihood and prior are both Gaussian, Thompson Sampling admits a closed-form posterior update, so we have a convenient ground-truth baseline for our approximate sampling methods.
@@
Real-World Tasks. Following Riquelme et al. (2018), we evaluate our MCMC Thompson Sampling algorithms on UCI datasets—Adult, Shuttle, MagicTelescope, Mushroom, Jester, Covertype, and RestaurantRatings—plus the Financial dataset and two vision benchmarks, MNIST_784 and CIFAR-10 (Krizhevsky, 2009; Elmachtoub et al., 2017). These are standard datasets studied in contextual bandits and multi-agent settings (Riquelme et al., 2018; Chaudhari et al., 2024), as they span a variety of features such as reward stochasticity, size, and number of optimal actions. Following the protocol of Riquelme et al. (2018); Kveton et al. (2020), each $N$-class example $x\!\in\!\mathbb{R}^{d}$ is converted into $K$ arm–context vectors $\smash{x^{(1)},\dots,x^{(k)}}$, where $x^{(i)}=x\cdot e_{i}$ where $e_{i}$ is the $i$’th standard basis vector of $\mathbb{R}^{d}$. Pulling arm $j$ returns reward $1$ if $j$ matches the ground-truth label, and 0 otherwise. We provide statistics of the benchmark datasets in Appendix F.
Network Architectures and Training Protocol. We adopt the training protocol and hyperparameter search strategy of Xu et al. (2022b): each method chooses the better of a fully connected two-layer MLP (width $100$) and four-layer MLP (width $50$); the activation is selected from {ReLU, LeakyReLU}. Networks are updated for $100$ SGD steps every round. FG-augmented variants inherit identical network architectures, optimizers, and training schedules as their non-FG counterparts to ensure fair comparisons. Unless stated otherwise, we run each bandit for T=10, [^0] T=10{,}000 rounds (Covertype: T=15, [^0] T=15{,}000). Additional implementation minutiae appear in Appendix A of Riquelme et al. (2018).
@@ 5 Results
Linear contextual bandits. Table 1 (first 3 columns) reports the final cumulative regret (mean $\pm$ sample standard deviation over 10 seeds) in the synthetic linear bandit setting. Across both dimensions $d\in\{20,40\}$, MALATS consistently attains the lowest cumulative regret: <62<62 when $d=20$ (vs 62.6 for unadjusted LMCTS, 73.0 for LinUCB when $\beta=10^{3}$) and around 100.6 when $d=40$ (vs 129.1 for LMCTS, 126.3 for LinUCB). Adding a feel-good optimism bonus (FGMALATS and SFGMALATS) of $\lambda=0.5$ does not improve regret in these Gaussian-conjugate settings. However, for smaller values of $\lambda$, such as $\lambda=0.01$ in Table 4, the performance of SFGMALATS improves significantly to 56.2 $\pm$ 22.8, surpassing the other vanilla-TS variants. When $d=40$, PHMCTS incurs $162.2\pm 12.5$ and SVRGLMCTS incurs 19236.8 $\pm$ 15560.8, indicating that shifting the posterior due to preconditioning and variance-reduction distorts the quadratic curvature of the true posterior.
Logistic contextual bandits ($d=20$). Table 1 (fourth column) reports the final cumulative regret in the logistic bandit setting. Although the reward is Bernoulli with a logistic link, in moderate signal‐to‐noise regimes, the posterior over the linear parameters is very nearly Gaussian (the log‐likelihood is strongly convex and concentrates around its mode). The main benefit of HMC to explore ‘curved’ [^3] This refers to the fact that the parameter $\theta$ lies in a Euclidean manifold on $\mathbb{R}^{d}$, rather than an affine subspace. or multimodal posteriors more effectively is not useful here, since we observe that simpler samplers like LMC and MALA closely match the closed-form Gaussian baselines. As expected, the exact‐Gaussian methods—LinUCB ($176.9\pm 41.9$) and LinTS ($179.9\pm 53.2$)—lead overall. Among the MCMC‐based samplers, FGLMCTS comes closest at $184.8\pm 38.0$, effectively matching LinTS with a smaller variance, while MALATS performs worse at $194.0\pm 76.9$.
Neural contextual bandits. Table 2 reports the final cumulative regret of a subset of our algorithms on several real-world problems, with complete results deferred to Section E.5. We additionally assess policy quality via simple regret in Table 3. When the posterior is approximated by SFG-TS, we observe a partial reversal: LMCTS becomes competitive with, and often outperforms, its feel-good extensions on the majority of the datasets. The only consistent advantage of the feel-good bonus emerges on MagicTelescope, where SFGLMCTS reduces cumulative regret by 27 and achieves the lowest simple regret of $82.8\pm 8.9$. Meanwhile, Neural-$\epsilon$-Greedy and NeuralUCB remain competitive, attaining the lowest regret on Mushroom and Adult, respectively. Overall, the greedy neural methods yield the lowest simple regret on three tasks.
@@ 5.1 Ablation Studies
(f) MNIST_784
Figure 1: Mean regret comparison on simulated bandit problems. The shaded band around each mean curve represents $\pm 1$ sample standard deviation across 5 independent runs.
Table 4: Final cumulative regret for synthetic datasets. Sweeps over feel-good parameter $\lambda$ at $\beta=10^{3}$.
@@ 6 Discussion
FG in Neural Contextual Bandits. Despite its stronger theoretical guarantees in the linear setting, FG-TS degrades rapidly in neural contextual bandits, where vanilla-TS remains more competitive. Stochastic gradients already inject substantial “free” exploration, and the FG bonus can push the sampler further from high-density regions. Neural posteriors are very non-Gaussian: mode-connectivity results show that SGD minima are linked by low-loss paths, forming a connected manifold rather than isolated valleys (Garipov et al., 2018; Draxler et al., 2019). Such topologies defies Gaussian or mixture assumptions of well-separated modes. Moreover, the quadratic surrogates underlying FG assume symmetric curvature, while Izmailov et al. (2019) show that SGD converges near the edge of wide, anisotropic basins with asymmetric directions, making quadratic fits reliable only locally. Consequently, FG-induced optimism amplifies model mismatch and yields sub-optimal arm choices. Vanilla LMCTS thus remains a strong baseline. In our experiments, moderate exploration, whether through SGD noise or $\epsilon$-greedy perturbations, outperforms the optimism of FG-TS. From a different perspective, our observations that FG-TS underperforms in the neural setting is the empirical consequence of applying an algorithm based on quadratic assumptions to this complex, non-quadratic setting.
Preconditioning. Our preconditioned samplers perform Langevin updates in a whitened parameter space injecting Gaussian noise with covariance proportional to the inverse design matrix $\mathbf{V}_{t}^{-1}$. When $\mathbf{V}_{t}$ is ill-conditioned, as it typically is in high dimensions, some directions receive huge stochastic kicks while others hardly move, leading to erratic exploration and exploding variance, as visible in SVRGLMCTS’s large regret. These imply a necessity to regularize or clip the spectrum of $\mathbf{V}_{t}^{-1}$ to keep noise injection balanced across directions, and to also include a Metropolis filter when using a mass-matrix preconditioner. Finally, our results on PSFGHMCTS indicate that it has a higher standard deviation than mean. This skew serves as an argument for incorporating additional metrics (such as the median) for studying the aggregate performance of MCMC contextual bandit algorithms.
@@
Damping. Table 31 reports LMC results with varying damping in the linear contextual bandit setting. Overdamped (vanilla) Langevin methods consistently outperform their underdamped counterparts, with the performance gap shrinking monotonically as the damping coefficient $\gamma$ increases. Although underdamped LMC enjoys stronger theoretical guarantees, its empirical performance deteriorates rapidly, which is likely due to posterior mismatch effects that are amplified at lower damping.
Run-to-run sensitivity. The variability in our experiments arises due to reward stochasticity, noise in mini-batch gradients within the MCMC samplers, and errors in posterior approximation with finitely many MCMC steps. Unfortunately, the contextual bandit feedback loop may amplify these errors, which is an important consideration for practitioners in the community. We reveal this sensitivity as a data point for the community to note that while MCMC methods are powerful, their application in online learning requires careful management of approximation quality for stability. Such variability in vanilla TS has already been noted (Mazumdar et al., 2020; Guha and Munagala, 2014); however, to our knowledge, our paper is the first to systematically document and analyze this for FGTS.